# A Bias–Variance Decomposition for Ensembles over Multiple Synthetic Datasets

### 06 Feb 2024 v1

@@ B.1 Background: Bregman Divergences
where $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is a strictly convex differentiable function. Many common error metrics, like MSE and cross entropy, can be expressed as expected values of a Bregman divergence. In fact, proper scoring rules666 Proper scoring rules are error metrics that are minimised by predicting the correct probabilities. can be characterised via Bregman divergences (Gneiting and Raftery, 2007; Kimpara et al., 2023). Table 1 shows how the metrics we consider are expressed as Bregman divergences (Gupta et al., 2022).
@@
Table 1: Common error metrics as Bregman divergences. $g$ denotes a prediction in regression and $p$ denotes predicted class probabilities in classification. $g^{(j)}$ and $p^{(j)}$ denote the predictions of different ensemble members. $y$ is the correct value in regression, and a one-hot encoding of the correct class in classification. The binary classification Brier score only looks at probabilities for one class. If the multiclass Brier score is used with two classes, it is twice the binary Brier score.
@@
B.2 Bregman Divergence Decomposition for Synthetic Data
We extend the Bregman divergence decomposition for ensembles from Gupta et al. (2022) to generative ensembles. To prove Theorem B.2, we use the following lemma. {lemma}[Gupta et al. 2022, Proposition 5.3] Let $X_{1},\dotsc,X_{m}$ be i.i.d.  random variables and let $\hat{X}=(\sum_{i=1}^{m}X_{i}^{*})^{*}$ be their dual average. Then $\mathcal{E}\hat{X}=\mathcal{E}X$, $\mathcal{V}\hat{X}\leq\mathcal{V}X$ and for any independent $Y$, $D(\operatorname*{\mathbb{E}}Y,\mathcal{E}\hat{X})=D(\operatorname*{\mathbb{E}}Y,\mathcal{E}X)$.
@@ B.2 Bregman Divergence Decomposition for Synthetic Data
Applying the generalised laws of expectation and variance, and Lemma B.2 to the variance term, we obtain:
@@
Appendix C Experimental Details
C.1 Datasets
In our experiments, we use 7 tabular datasets. For four of them, the downstream prediction task is regression, and for the other three, the prediction task is binary classification. Table 2 lists some general information on the datasets. We use 25% of the real data as a test set, with the remaining 75% being used to generate the synthetic data, for all of the datasets. All experiments are repeated several times, with different train-test splits for each repeat.
Table 2: Details on the datasets used in the experiments. # Cat. and # Num. are the numbers of categorical and numerical features, not counting the target variable. For datasets with removed rows, the table shows the number of rows after the removals.
@@ C.1 Datasets
(Nash et al., 1995) The abalone dataset contains information on abalones, with the task of predicting the number of rings on the abalone from the other information.
@@
(https://www.census.gov/programs-surveys/acs/microdata/documentation.2018.html) This dataset contains several variables from the American community survey (ACS) of 2018, with the task of predicting a person’s income from the other features. Specifically, the variables we selected are AGEP (age), COW (employer type), SCHL (education), MAR (marital status), WKHP (working hours), SEX, RAC1P (race), and the target PINCP (income). We take a subset of 50000 datapoints from the California data, and log-transform the target variable. We used the folktables package (Ding et al., 2021) to download the given subset of the data.
@@
(Kohavi and Becker, 1996) The UCI Adult dataset contains general information on people, with the task of predicting whether their income is over \$50000. We drop rows with any missing values.
@@
(Wolberg et al., 1995) The breast cancer dataset contains features derived from images of potential tumors, with the task of predicting whether the potential tumor is benign or malignant.
@@
(https://scikit-learn.org/stable/datasets/real_world.html#california-housing-dataset) The california housing dataset contains information on housing districts, specifically census block groups, in California. The task is predicting the median house value in the district. We removed outlier rows where the average number of rooms is at least 50, or the average occupancy is at least 30. According to the dataset description, these likely correspond to districts with many empty houses. We log-transformed the target variable, as well as the population and median income features.
@@
(Hofmann, 1994) The German credit dataset contains information on a bank’s customers, with the task of predicting whether the customers are “good” or “bad”.
@@
(https://www.kaggle.com/datasets/mirichoi0218/insurance/data) The insurance dataset contains general information on people, like age, gender and BMI, as well as the amount they charged their medical insurance, which is the variable to predict. We take a log transform of the target variable before generating synthetic data.
C.2 Downstream Prediction Algorithms
@@
We use the scikit-learn [^7] https://scikit-learn.org/stable/index.html implementations of all of the downstream algorithms, which includes probability predictions for all algorithms on the classification tasks. We standardise the data before training for all downstream algorithms except the tree-based algorithms, specifically decision tree, random forest, and gradient boosted trees. This standardisation is done just before downstream training, so the input to the synthetic data generation algorithms is not standardised. We use the default hyperparameters of scikit-learn for all downstream algorithms except MLP, where we increased the maximum number of iterations to 1000, as the default was not enough to converge on some datasets. In particular, this means that decision trees are trained to interpolate the training data, resulting in high variance of the predictions.
Appendix D Extra Results
@@
Figure 5: All error metrics on the Adult dataset. Note the logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure 6: All error metrics on the breast cancer dataset. Note the he logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure 7: All error metrics on the German credit dataset. Note the he logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure 8: MSE prediction on the first two regression datasets. The linear regression measured MSE line for synthpop with ACS 2018 data is cut off due to excluding repeats with extremely large MSE ($\geq 10^{6}$). 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure 9: MSE prediction on the last two regression datasets. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure 10: Brier score prediction on two classification datasets. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
Figure 11: Brier score prediction on German credit data. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure 12: Random forest performance prediction on the regression datasets in (a) and classification datasets in (b). The prediction is reasonably accurate on the datasets with accurate estimates of the error. On the other datasets, the prediction can have high variance. The lines show averages over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootsrapping over the repeats.
Table 3: Table of synthetic data generator comparison results from Figure 1. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
@@
Table 4: Table of results from the Abalone dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
@@
Table 5: Table of results from the ACS 2018 dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats. The nans for linear regression are caused by excluding repeats with extremely large MSE ($\geq 10^{6}$).
@@
Table 6: Table of results from the California Housing dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
@@
Table 7: Table of results from the Insurance dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
@@
Table 8: Table of results from the Adult dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
@@
Table 9: Table of results from the Breast Cancer dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
@@
Table 10: Table of results from the German Credit dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
@@
Table 11: Table of results from predicting MSE on the ACS 2018 dataset. The numbers are the mean measured and predicted MSEs $\pm$ standard deviations from 3 repeats. The nans for linear regression are due to excluding some repeats that had an extremely large MSE ($\geq 10^{6}$).

### 23 May 2024 v2

@@ C.1 Background: Bregman Divergences
where $F\colon\mathbb{R}^{d}\to\mathbb{R}$ is a strictly convex differentiable function. Many common error metrics, like MSE and cross entropy, can be expressed as expected values of a Bregman divergence. In fact, proper scoring rules555 Proper scoring rules are error metrics that are minimised by predicting the correct probabilities. can be characterised via Bregman divergences (Gneiting & Raftery, 2007; Kimpara et al., 2023). Table S1 shows how the metrics we consider are expressed as Bregman divergences (Gupta et al., 2022).
@@
Table S1: Common error metrics as Bregman divergences. $g$ denotes a prediction in regression and $p$ denotes predicted class probabilities in classification. $g^{(j)}$ and $p^{(j)}$ denote the predictions of different ensemble members. $y$ is the correct value in regression, and a one-hot encoding of the correct class in classification. The binary classification Brier score only looks at probabilities for one class. If the multiclass Brier score is used with two classes, it is twice the binary Brier score.
@@
C.2 Bregman Divergence Decomposition for Synthetic Data
We extend the Bregman divergence decomposition for ensembles from Gupta et al. (2022) to generative ensembles. To prove Theorem C.2, we use the following lemma. {lemma}[Gupta et al. 2022, Proposition 5.3] Let $X_{1},\dotsc,X_{m}$ be i.i.d.  random variables and let $\hat{X}=(\sum_{i=1}^{m}X_{i}^{*})^{*}$ be their dual average. Then $\mathcal{E}\hat{X}=\mathcal{E}X$, $\mathcal{V}\hat{X}\leq\mathcal{V}X$ and for any independent $Y$, $D(\operatorname*{\mathbb{E}}Y,\mathcal{E}\hat{X})=D(\operatorname*{\mathbb{E}}Y,\mathcal{E}X)$.
@@ C.2 Bregman Divergence Decomposition for Synthetic Data
Applying the generalised laws of expectation and variance, and Lemma C.2 to the variance term, we obtain:
@@
Appendix D Experimental Details
D.1 Datasets
In our experiments, we use 7 tabular datasets. For four of them, the downstream prediction task is regression, and for the other three, the prediction task is binary classification. Table S2 lists some general information on the datasets. We use 25% of the real data as a test set, with the remaining 75% being used to generate the synthetic data, for all of the datasets. All experiments are repeated several times, with different train-test splits for each repeat.
Table S2: Details on the datasets used in the experiments. # Cat. and # Num. are the numbers of categorical and numerical features, not counting the target variable. For datasets with removed rows, the table shows the number of rows after the removals.
@@ D.1 Datasets
(Nash et al., 1995, CC BY 4.0) The abalone dataset contains information on abalones, with the task of predicting the number of rings on the abalone from the other information.
@@
(https://www.census.gov/programs-surveys/acs/microdata/documentation.2018.html, license: https://www.census.gov/data/developers/about/terms-of-service.html) This dataset contains several variables from the American community survey (ACS) of 2018, with the task of predicting a person’s income from the other features. Specifically, the variables we selected are AGEP (age), COW (employer type), SCHL (education), MAR (marital status), WKHP (working hours), SEX, RAC1P (race), and the target PINCP (income). We take a subset of 50000 datapoints from the California data, and log-transform the target variable. We used the folktables package (Ding et al., 2021) to download the given subset of the data.
@@
(Kohavi & Becker, 1996, CC BY 4.0) The UCI Adult dataset contains general information on people, with the task of predicting whether their income is over \$50000. We drop rows with any missing values.
Reduced Adult
(Kohavi & Becker, 1996, CC BY 4.0) This dataset is the UCI Adult dataset with a reduced set of features. The subset of features is age, workclass, education, marital-status, race, gender, capital-gain, capital-loss and hours-per-week. [^6] This subset was used in the NAPSU-MQ experiments of Räisä et al. (2023b) We binarise capital-gain and capital-loss to indicate whether the original value is positive or not for both synthetic data generation and downstream prediction. We discretise age and hours-per-week to 5 categories for synthetic data generation, and converted back to continuous values for the downstream prediction, as our synthetic data generators require discrete data. We only remove rows with missing values in the included columns, so the number of rows is larger with the reduced set of features.
@@
(Wolberg et al., 1995, CC BY 4.0) The breast cancer dataset contains features derived from images of potential tumors, with the task of predicting whether the potential tumor is benign or malignant.
@@
(https://scikit-learn.org/stable/datasets/real_world.html#california-housing-dataset, license not mentioned) The california housing dataset contains information on housing districts, specifically census block groups, in California. The task is predicting the median house value in the district. We removed outlier rows where the average number of rooms is at least 50, or the average occupancy is at least 30. According to the dataset description, these likely correspond to districts with many empty houses. We log-transformed the target variable, as well as the population and median income features.
@@
(Hofmann, 1994, CC BY 4.0) The German credit dataset contains information on a bank’s customers, with the task of predicting whether the customers are “good” or “bad”.
@@
(https://www.kaggle.com/datasets/mirichoi0218/insurance/data, Database Contents License (DbCL) v1.0) The insurance dataset contains general information on people, like age, gender and BMI, as well as the amount they charged their medical insurance, which is the variable to predict. We take a log transform of the target variable before generating synthetic data.
D.2 Downstream Prediction Algorithms
@@
We use the scikit-learn [^7] https://scikit-learn.org/stable/index.html, BSD 3-Clause License implementations of all of the downstream algorithms, which includes probability predictions for all algorithms on the classification tasks. We standardise the data before training for all downstream algorithms except the tree-based algorithms, specifically decision tree, random forest, and gradient boosted trees. This standardisation is done just before downstream training, so the input to the synthetic data generation algorithms is not standardised. We use the default hyperparameters of scikit-learn for all downstream algorithms except MLP, where we increased the maximum number of iterations to 1000, as the default was not enough to converge on some datasets. In particular, this means that decision trees are trained to interpolate the training data, resulting in high variance of the predictions.
D.3 DP Experiment
Both AIM (McKenna et al., 2022) and NAPSU-MQ (Räisä et al., 2023b) generate synthetic data based on noisy values of marginal queries on the real data, which count how many rows of the data have given values for given variables. AIM includes a mechanism that chooses a subset of queries to measure under DP from a potentially very large workload. We set the workload to be all marginal queries over two variables. NAPSU-MQ does not include a query selection mechanism, so we first run AIM with $\epsilon=0.5$, $\delta=\frac{1}{2}n^{-2}$ to select a subset of marginals and then run NAPSU-MQ with the selected marginals with $\epsilon=1$, $\delta=\frac{1}{2}n^{-2}$, which results in the same privacy bounds ($\epsilon=1.5$, $\delta=n^{-2}$; Dwork & Roth, 2014) that we used for AIM. We split the privacy budget between the $m$ synthetic datasets in AIM by dividing the zero-concentrated DP (Bun & Steinke, 2016) parameter that AIM uses internally by $m$, which keeps the total privacy budget fixed.
[…]
D.4 Computational Resources
We ran the synthetic data generation algorithms DDPM, TVAE and CTGAN on a single GPU, synthpop, AIM and NAPSU-MQ on CPU, and all downstream analysis on CPU, all in a cluster environment.
Appendix E Extra Results
E.1 Synthetic Data Generation Algorithms
We compare several synthetic data generation algorithms to see which algorithms are most interesting for subsequent experiments. We use the California housing dataset, where the downstream task is regression. The algorithms we compare are DDPM (Kotelnikov et al., 2023), TVAE (Xu et al., 2019), CTGAN (Xu et al., 2019) and synthpop (Nowok et al., 2016). DDPM, TVAE and CTGAN are a diffusion model, a variational autoencoder and a GAN that are designed for tabular data. We use the implementations from the synthcity library [^11] https://github.com/vanderschaarlab/synthcity, Apache-2.0 license for these. Synthpop generates synthetic data by sampling one column from the real data, and generating the other columns by sequentially training a predictive model on the real data, and predicting the next column from the already generated ones. We use the implementation from the authors (Nowok et al., 2016). [^12] License: GPL-3
The results of the comparison are shown in Figure S1. We see that synthpop and DDPM with MSE loss generally outperform the other generation algorithms, so we select them for the subsequent experiments. There is very little difference between the two variants of synthpop, so we choose the “proper” variant due to its connection with the Bayesian reasoning for using multiple synthetic datasets.
Figure S1: Comparison of synthetic data generation algorithms for several prediction algorithms on the California housing dataset, with 1 to 10 synthetic datasets. DDPM and synthpop achieve smaller MSE in the downstream predictions, so they were selected for further experiments. SP-P and SP-IP are the proper and improper variants of synthpop, and DDPM-KL is DDPM with KL divergence loss. 1-NN and 5-NN are nearest neighbours with 1 and 5 neighbours. The dashed black lines show the performance of each prediction algorithm on the real data, and the solid black line shows the performance of the best predictor, random forest, on the real data. The results are averaged over 3 repeats, with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. Linear regression was omitted, as it had nearly identical results as ridge regression. Table S3 in the Appendix contains the numbers in the plots, including ridge regression.
E.2 Extra Plots
@@
Figure S2: MSE on regression datasets of the ensemble of downstream predictors, with varying number of synthetic datasets $m$ from DDPM. Increasing the number of synthetic datasets generally decreases MSE, especially for decision trees and 1-NN. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and ridge regression (RR). The black line is the MSE of the best predictor on real data. The results are averaged over 3 repeats. The error bars are 95% confidence intervals formed by bootstrapping over the repeats. We omitted linear regression from the plots, as it had almost identical results to ridge regression. Tables S4 to S7 contain the numbers from the plots, including linear regression.
@@
Figure S3: All error metrics on the Adult dataset. Note the logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure S4: All error metrics on the breast cancer dataset. Note the he logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
Figure S5: All error metrics on the German credit dataset. Note the he logarithmic scale on the cross entropy y-axis. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure S6: MSE prediction on the first two regression datasets. The linear regression measured MSE line for synthpop with ACS 2018 data is cut off due to excluding repeats with extremely large MSE ($\geq 10^{6}$). 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure S7: MSE prediction on the last two regression datasets. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure S8: Brier score prediction on two classification datasets. 1-NN and 5-NN are nearest neighbours with 1 or 5 neighbours. The results are averaged over 3 repeats with different train-test splits. The error bands are 95% confidence intervals formed by bootstrapping over the repeats.
@@
Figure S9: Random forest performance prediction on the regression datasets in (a) and classification datasets in (b). The prediction is reasonably accurate on the datasets with accurate estimates of the error. On the other datasets, the prediction can have high variance. The lines show averages over 3 different train-test splits and 3 repeats of model training per split. The error bands are 95% confidence intervals formed by bootsrapping over the repeats and different splits.
Figure S10: All error metrics on the Adult dataset with reduced features and DP synthetic data generation. Note the he logarithmic scale on the cross entropy y-axis. The privacy parameters are $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$. The predictors are nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), decision tree (DT), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM) and logistic regression (LogR). The black line show the loss of the best downstream predictor trained on real data. The results are averaged over 3 repeats with different train-test splits. The error bars are 95% confidence intervals formed by bootstrapping over the repeats.
E.3 Result Tables
Table S3: Table of synthetic data generator comparison results from Figure S1. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
@@ E.3 Result Tables
Table S4: Table of results from the Abalone dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
@@
Table S5: Table of results from the ACS 2018 dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats. The nans for linear regression are caused by excluding repeats with extremely large MSE ($\geq 10^{6}$).
@@
Table S6: Table of results from the California Housing dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
@@
Table S7: Table of results from the Insurance dataset. The numbers are the mean MSE $\pm$ standard deviation from 3 repeats.
@@
Table S8: Table of results from the Adult dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
@@
Table S9: Table of results from the Breast Cancer dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
@@
Table S10: Table of results from the German Credit dataset. The numbers are the mean Brier score $\pm$ standard deviation from 3 repeats.
@@
Table S11: Table of results from predicting MSE on the ACS 2018 dataset. The numbers are the mean measured and predicted MSEs $\pm$ standard deviations from 3 repeats. The nans for linear regression are due to excluding some repeats that had an extremely large MSE ($\geq 10^{6}$).